# Nonuniqueness of Leray–Hopf solutions to the unforced incompressible 3D Navier–Stokes Equation

### v1

@@ 4.5 Numerical results
Although comprehensive numerical details are presented in subsequent sections, we summarize the key results here for convenience. We will conclude the propositions by verifying Assumption 6 and 7. We remark that our verifications are carried effectively in a subspace of $H^{1}_{\sigma}$, leveraging axissymmetry and parities, to reduce the number of bases in our finite-rank approximation.
@@ 4.5.1 Verification of Proposition 1
$M_{4}^{U}=(\frac{1}{4}-\eta_{1}-M_{2}^{U})^{2}-4M_{3}^{U}\varepsilon^{U}>0,$
where the residual satisfies
$\varepsilon^{U}\leqslant 7\times 10^{-7}.$
For $\eta_{1}$, we take $\eta_{1}=0.191$ and verify that
$\frac{\nabla\overline{U}+\nabla^{T}\overline{U}}{2}+\eta_{1}I+Q\succeq 0.$
Then $M_{4}^{U}$ reduces to
$M_{4}^{U}\geqslant(0.059-M_{2}^{U})^{2}-2.8\times 10^{-6}\times M_{3}^{U}.$
@@
$M_{1}^{U}=K^{U}_{5}+\frac{K_{6}^{U}}{(1-\varepsilon^{U}_{4})^{2}},\quad K^{U}_{5}=\|\lambda_{\max}(Q)\|_{L^{\infty}}\sup_{u\neq 0,u\bot_{L^{2}}V^{U}_{N}}\frac{\|u\|^{2}_{L^{2}(\Omega)}}{\|u\|^{2}_{E}},\quad$
@@
$K_{6}^{U}=\sup_{u\in V_{N}^{U},u\neq 0}\frac{\langle u,Qu\rangle_{\Omega}-\sum_{j=1}^{m}\overline{\lambda}_{j}^{-1}\langle u,Q\overline{\psi}_{j}^{U}\rangle_{\Omega}^{2}}{\|u\|^{2}_{E}}\leqslant 0.46.$
We choose $N=1257$, $m=19$ and get:
$\|\lambda_{\max}(Q)\|_{\infty}\leqslant 4.5,\quad\sup_{u\neq 0,u\bot_{L^{2}}V^{U}_{N}}\frac{\|u\|^{2}_{L^{2}(\Omega)}}{\|u\|^{2}_{E}}\leqslant 0.1024.$
It follows that
$K^{U}_{5}\leqslant 0.47.$
$\varepsilon^{U}_{3}$ and $\varepsilon^{U}_{4}$ arise from machine precision and are bounded by
$\varepsilon^{U}_{3},\varepsilon^{U}_{4}\leqslant 10^{-10}.$
@@
$M_{1}^{U}\leqslant 0.47+\frac{0.46}{(1-10^{-10})^{2}}\leqslant 0.9301.$
The constant $K_{7}^{U}$ defined by
$K_{7}^{U}=\sqrt{\operatorname{tr}(G_{Q^{2}}G_{r})}\leqslant 0.0002.$
Therefore,
$M_{2}^{U}\leqslant 0.005\times 0.9301+10^{-10}\times 4.5+0.0002\leqslant 0.005.$
@@
$M_{3}^{U}$ is defined by
$M_{3}^{U}=\sqrt{\|G^{U}_{Q^{2}}\|\sum_{j=1}^{m}\|\nabla\overline{\phi}_{j}^{U}\|_{L^{\infty}}^{2}}+\|G^{U}_{Q^{2}}\|\sqrt{\sum_{j=1}^{m}\sum_{i=1}^{m}\|\overline{\phi}_{i}^{U}\cdot\nabla\overline{\phi}_{j}^{U}\|_{L^{2}}^{2}}.$
Numerical results show that
$\sqrt{\sum_{j=1}^{m}\|\nabla\overline{\phi}_{j}^{U}\|_{L^{\infty}}^{2}}\leqslant 0.9,\quad\sqrt{\sum_{j=1}^{m}\sum_{i=1}^{m}\|\overline{\phi}_{i}^{U}\cdot\nabla\overline{\phi}_{j}^{U}\|_{L^{2}}^{2}}\leqslant 0.11,\quad\|G^{U}_{Q^{2}}\|\leqslant 0.025,$
which lead to,
$M_{3}^{U}\leqslant\sqrt{0.025}\times 0.9+0.025\times 0.11\leqslant 0.16.$
@@
$M_{4}^{U}\geqslant(\frac{1}{4}-0.191-0.005)^{2}-4\times 0.16\times 7\times 10^{-7}\geqslant 0.0029.$
@@
$x_{0}^{U}=\frac{2\varepsilon^{U}}{\left(\frac{1}{4}-\eta_{1}-M_{2}^{U}+\sqrt{M_{4}^{U}}\right)}\leqslant\frac{2\times 7\times 10^{-7}}{0.25-0.191-0.005+\sqrt{0.0029}}\leqslant 1.3\times 10^{-5},$
@@
$x_{1}^{U}=\sqrt{\frac{\frac{1}{4}-\eta_{1}-M_{2}^{U}}{1-M_{1}^{U}}}\frac{x_{0}^{U}}{2}\leqslant\sqrt{\frac{\frac{1}{4}-0.191}{1-0.9301}}\times\frac{1.3\times 10^{-5}}{2}\leqslant 6\times 10^{-6}.$
@@ 4.5.2 Verification of Proposition 2
$M_{4}^{v}=(M_{5}^{v})^{2}-4M_{3}^{v}\varepsilon^{v,\operatorname{tot}}>0,$
where we denote
$\varepsilon^{v,\operatorname{tot}}$ $:=$ $\varepsilon^{v}+\|U\|_{L^{2}}\|\nabla\overline{v}\|_{L^{\infty}}+\|\nabla U\|_{L^{2}}\|\overline{v}\|_{L^{\infty}},$
$M_{5}^{v}$ $:=$ $\frac{1}{4}-\eta_{1}-\overline{\lambda}-2K_{4}\|\nabla U\|_{L^{2}}-|\lambda^{\operatorname{con}}|-M_{2}^{v}.$
The eigenvalue $\overline{\lambda}$ is
$\overline{\lambda}\approx-0.11314203274385938\leqslant-0.113.$
Step 1: Estimate of $\varepsilon^{v,\operatorname{tot}}$ and $|\lambda^{\operatorname{con}}|$.
The residual satisfies
$\varepsilon^{v}\leqslant 1.8\times 10^{-5}.$
The $L^{\infty}$ norm of $\overline{v}$ and $\nabla\overline{v}$ are:
$\|\nabla\overline{v}\|_{L^{\infty}}\leqslant 3,\|\overline{v}\|_{L^{\infty}}\leqslant 2.3.$
Therefore, $|\lambda^{\operatorname{con}}|$ and $\varepsilon^{v,\operatorname{tot}}$ are bounded by
$|\lambda^{\operatorname{con}}|$ $\leqslant$ $\|\nabla U\|_{L^{2}}\|\overline{v}\|_{L^{\infty}}\leqslant 6\times 10^{-6}\times 2.3\leqslant 1.4\times 10^{-5},$
$\varepsilon^{v,\operatorname{tot}}$ $\leqslant$ $1.8\times 10^{-5}+1.3\times 10^{-5}\times 3+6\times 10^{-6}\times 2.3\leqslant 7.1\times 10^{-5}.$
Step 2: Estimate of $M_{5}^{v}$.
$K_{4}\leqslant 0.091$ is introduced in Lemma 6. We can simplify $M_{5}^{v}$ into
$M_{5}^{v}\geqslant\frac{1}{4}-0.191+0.113-2\times 0.091\times 6\times 10^{-6}-1.4\times 10^{-5}-M_{2}^{v}\geqslant 0.17-M_{2}^{v}.$
@@
We choose $N=1260,m=24$ and get:
$M_{1}^{v}$ $=$ $K^{v}_{5}+\frac{K_{6}^{v}}{(1-\varepsilon^{v}_{4})^{2}},$
$K_{7}^{v}$ $=$ $\sqrt{\operatorname{tr}(G_{Q^{2}}^{v}G_{r}^{v})}\leqslant 0.00027,$
$K_{6}^{v}$ $=$ $\sup_{u\in V_{N}^{v},u\neq 0}\frac{\langle u,Qu\rangle_{\Omega}-\sum_{j=1}^{m}\overline{\lambda}_{j}^{-1}\langle u,Q\overline{\psi}_{j}^{v}\rangle_{\Omega}^{2}}{\|u\|^{2}_{E}}\leqslant 0.4532.$
Therefore, $M_{1}^{v}$, $M_{2}^{v}$ are bounded by
$M_{1}^{v}$ $\leqslant$ $0.47+\frac{0.4532}{(1-10^{-10})^{2}}\leqslant 0.924,$
$M_{2}^{v}$ $\leqslant$ $0.005\times 0.924+4.5\times 10^{-10}+0.00027\leqslant 0.0049.$
$M_{5}^{v}$ thus has the following lower bound:
$M_{5}^{v}\geqslant 0.17-0.0049=0.1651.$
@@
$M_{0}=\|P_{\overline{v}}(\frac{1}{2}x\cdot\nabla\overline{v}-\Delta\overline{v}-\overline{U}\cdot\nabla\overline{v}+\overline{v}\cdot\nabla^{T}\overline{U})\|_{L^{2}}+\|U\|_{L^{2}}\|\nabla\overline{v}\|_{L^{\infty}}+\|\nabla U\|_{L^{2}}\|\overline{v}\|_{L^{\infty}}.$
We collect the numerical computations:
$\|P_{\overline{v}}(\frac{1}{2}x\cdot\nabla\overline{v}-\Delta\overline{v}-\overline{U}\cdot\nabla\overline{v}+\overline{v}\cdot\nabla^{T}\overline{U})\|_{L^{2}}\leqslant 6.3581,\quad\sqrt{\|G_{Q^{2}}^{v}\|\sum_{j=1}^{m}\|\overline{\phi}_{j}^{v}\|_{L^{2}}^{2}}\leqslant 0.03,$
which yield the following approximate values:
$M_{0}$ $\leqslant$ $6.3581+1.3\times 10^{-5}\times 3+6\times 10^{-6}\times 2.3\leqslant 6.36$
$M_{3}^{v}$ $\leqslant$ $6.36\times(1+0.03)^{2}\leqslant 7,$
$M_{4}^{v}$ $\geqslant$ $0.1651^{2}-4\times 7\times 7.1\times 10^{-5}\geqslant 0.025.$
@@
$x_{0}^{v}$ $=$ $\frac{2\varepsilon^{v,\operatorname{tot}}}{M_{5}^{v}+\sqrt{M_{4}^{v}}}\leqslant\frac{2\times 7.1\times 10^{-5}}{0.1651+\sqrt{0.025}}\leqslant 4.4\times 10^{-4},$
$x_{1}^{v}$ $=$ $\sqrt{\frac{M_{5}^{v}}{1-2K_{4}\|\nabla U\|_{L^{2}}-M_{1}^{v}}}\frac{x_{0}^{v}}{2}\leqslant\sqrt{\frac{0.1651}{1-2\times 0.091\times 6\times 10^{-6}-0.924}}\times\frac{4.4\times 10^{-4}}{2}\leqslant 3.3\times 10^{-4}.$
@@
$|\lambda^{\operatorname{lin}}(v)|\leqslant M_{0}\|v\|_{L^{2}}\leqslant M_{0}(1+\sqrt{\sum_{j=1}^{m}\|G_{Q^{2}}^{v}\|\|\overline{\phi}_{j}^{v}\|_{L^{2}}^{2}})x_{0}^{v}\leqslant 6.36\times(1+0.03)\times 4.4\times 10^{-4}\leqslant 0.00293.$
@@
$|\lambda|\leqslant|\lambda^{\operatorname{lin}}(v)|+|\lambda^{\operatorname{con}}|\leqslant 0.003\ll|\overline{\lambda}|\approx 0.113.$
This concludes the numerical verification and thereby completes the proof of Proposition 2.

### v2

@@ 4.5 Numerical results
Although comprehensive numerical details are presented in subsequent sections, we summarize the key results here for convenience. In particular, we complete the proofs of the propositions by verifying Assumptions 6 and 7. We remark that our verifications are carried out effectively in a subspace of $H^{1}_{\sigma}$, leveraging axissymmetry and parities, to reduce the number of basis functions in our finite-rank approximation. We begin by listing all quantities verified by numerics:
$\|\mathrm{E}^{U}\|_{L^{2}}$ $\leqslant$ $\varepsilon^{U}=6.5\times 10^{-7},$ (4.44)
$\frac{\nabla\bar{U}+\nabla^{T}\bar{U}}{2}+\eta_{1}I+Q$ $\succeq$ $0,\quad\eta_{1}=0.199,$ (4.45)
$\|\lambda_{\max}(Q)\|_{L^{\infty}}$ $\leqslant$ $4.5,$ (4.46)
$\|\mathrm{E}^{v}\|_{L^{2}}$ $\leqslant$ $\varepsilon^{v}=1.9\times 10^{-5},$ (4.47)
$\overline{\lambda}$ $\approx$ $-0.11314203274385938\leqslant-0.113,$ (4.48)
$\|\overline{v}\|_{L^{\infty}}$ $\leqslant$ $2.52,$ (4.49)
$\|\nabla\overline{v}\|_{L^{\infty}}$ $\leqslant$ $3.4,$ (4.50)
$\|P_{\bar{v}}(\frac{1}{2}x\cdot\nabla\bar{v}-\Delta\bar{v}-\bar{U}\cdot\nabla\bar{v}+\bar{v}\cdot\nabla^{T}\bar{U})\|_{L^{2}}$ $\leqslant$ $6.38,$ (4.51)
$\sup_{u\neq 0,u\bot_{L^{2}}V^{U}_{N}}\frac{\|u\|^{2}_{L^{2}(\Omega)}}{\|u\|^{2}_{E}},\sup_{u\neq 0,u\bot_{L^{2}}V^{v}_{N}}\frac{\|u\|^{2}_{L^{2}(\Omega)}}{\|u\|^{2}_{E}}$ $\leqslant$ $0.1024,$ (4.52)
$\sup_{u\in V_{N}^{U},u\neq 0}\frac{\langle u,Qu\rangle_{\Omega}-\sum_{j=1}^{m}\overline{\lambda}_{j}^{-1}\langle u,Q\overline{\psi}_{j}^{U}\rangle_{\Omega}^{2}}{\|u\|^{2}_{E}}$ $\leqslant$ $0.46,$ (4.53)
$\sup_{u\in V_{N}^{v},u\neq 0}\frac{\langle u,Qu\rangle_{\Omega}-\sum_{j=1}^{m}\overline{\lambda}_{j}^{-1}\langle u,Q\overline{\psi}_{j}^{v}\rangle_{\Omega}^{2}}{\|u\|^{2}_{E}}$ $\leqslant$ $0.46,$ (4.54)
$\varepsilon^{U}_{3},\varepsilon^{U}_{4},\varepsilon^{v}_{3},\varepsilon^{v}_{4}$ $\leqslant$ $5.3\times 10^{-5},$ (4.55)
$\sqrt{\operatorname{tr}(G_{Q^{2}}^{U}G_{r}^{U})}$ $\leqslant$ $0.0012,$ (4.56)
$\|G^{U}_{Q^{2}}\|_{F}$ $\leqslant$ $0.025,$ (4.57)
$\sqrt{\sum_{j=1}^{m}\sum_{l=1}^{m}\|\overline{\phi}_{l}^{U}\cdot\nabla\overline{\phi}_{j}^{U}\|^{2}_{L^{2}}}$ $\leqslant$ $0.12,$ (4.58)
$\sqrt{\sum_{j=1}^{m}\|\nabla\overline{\phi}_{j}^{U}\|^{2}_{L^{\infty}}}$ $\leqslant$ $1.33,$ (4.59)
$\sqrt{\sum_{j=1}^{m}\|\overline{\phi}_{j}^{U}\|^{2}_{L^{2}}}$ $\leqslant$ $0.36,$ (4.60)
$\sqrt{\operatorname{tr}(G_{Q^{2}}^{v}G_{r}^{v})}$ $\leqslant$ $0.004,$ (4.61)
$\|G^{v}_{Q^{2}}\|_{F}$ $\leqslant$ $0.029,$ (4.62)
$\sqrt{\sum_{j=1}^{m}\|\overline{\phi}_{j}^{v}\|^{2}_{L^{2}}}$ $\leqslant$ $1.02\;.$ (4.63)
@@ 4.5.1 Verification of Proposition 1
$M_{4}^{U}=(\frac{1}{4}-\eta_{1}-M_{2}^{U})^{2}-4M_{3}^{U}\varepsilon^{U}>0.$
Since the residual satisfies
$\varepsilon^{U}\leqslant 6.5\times 10^{-7},$
and we take $\eta_{1}=0.199$, $M_{4}^{U}$ reduces to
$M_{4}^{U}\geqslant(0.051-M_{2}^{U})^{2}-2.8\times 10^{-6}\times M_{3}^{U}.$
@@
$M_{1}^{U}=K^{U}_{5}+\frac{K_{6}^{U}}{(1-\varepsilon^{U}_{4})^{2}},\quad$
@@
$K_{6}^{U}=\sup_{u\in V_{N}^{U},u\neq 0}\frac{\langle u,Qu\rangle_{\Omega}-\sum_{j=1}^{m}\overline{\lambda}_{j}^{-1}\langle u,Q\overline{\psi}_{j}^{U}\rangle_{\Omega}^{2}}{\|u\|^{2}_{E}}\leqslant 0.46,$
and
$K^{U}_{5}=\|\lambda_{\max}(Q)\|_{L^{\infty}}\sup_{u\neq 0,u\bot_{L^{2}}V^{U}_{N}}\frac{\|u\|^{2}_{L^{2}(\Omega)}}{\|u\|^{2}_{E}}\leqslant 0.47$
@@
$M_{1}^{U}\leqslant 0.47+\frac{0.46}{(1-5.3\times 10^{-5})^{2}}\leqslant 0.9301.$
By the definition of the constant $K_{7}^{U}$, we can bound it as follows:
$K_{7}^{U}=\sqrt{\operatorname{tr}(G^{U}_{Q^{2}}G^{U}_{r})}\leqslant 0.0012.$
Therefore, we obtain
$M_{2}^{U}\leqslant 0.005\times 0.9301+5.3\times 10^{-5}\times 4.5+0.0012\leqslant 0.0061.$
@@
$M_{3}^{U}$ is defined and bounded by
$M_{3}^{U}=\sqrt{\|G^{U}_{Q^{2}}\|\sum_{j=1}^{m}\|\nabla\overline{\phi}_{j}^{U}\|_{L^{\infty}}^{2}}+\|G^{U}_{Q^{2}}\|\sqrt{\sum_{j=1}^{m}\sum_{i=1}^{m}\|\overline{\phi}_{i}^{U}\cdot\nabla\overline{\phi}_{j}^{U}\|_{L^{2}}^{2}}\leqslant\sqrt{0.025}\times 1.33+0.025\times 0.12\leqslant 0.22.$
@@
$M_{4}^{U}\geqslant(\frac{1}{4}-0.199-0.0061)^{2}-4\times 0.22\times 6.5\times 10^{-7}\geqslant 0.002.$
@@
$x_{0}^{U}=\frac{2\varepsilon^{U}}{\left(\frac{1}{4}-\eta_{1}-M_{2}^{U}+\sqrt{M_{4}^{U}}\right)}\leqslant\frac{2\times 6.5\times 10^{-7}}{0.25-0.199-0.005+\sqrt{0.002}}\leqslant 1.44\times 10^{-5},$
@@
$x_{1}^{U}=\sqrt{\frac{\frac{1}{4}-\eta_{1}-M_{2}^{U}}{1-M_{1}^{U}}}\frac{x_{0}^{U}}{2}\leqslant\sqrt{\frac{\frac{1}{4}-0.199}{1-0.9301}}\times\frac{1.44\times 10^{-5}}{2}\leqslant 6.2\times 10^{-6}.$
By the estimates (4.29) and (4.30), $U_{2}$ is bounded by:
$y_{0}^{U}:=\|U_{2}\|_{L^{2}}$ $\leqslant$ $\|U_{1}\|_{L^{2}}\|G_{Q^{2}}^{U}\|_{F}\sqrt{\sum_{j=1}^{m}\|\overline{\phi}_{j}^{U}\|_{L^{2}}^{2}}\leqslant 1.44\times 10^{-5}\times 0.025\times 0.36\leqslant 1.3\times 10^{-7},$
$y_{1}^{U}:=\|\nabla U_{2}\|_{L^{\infty}}$ $\leqslant$ $\|U_{1}\|_{L^{2}}\|G_{Q^{2}}^{U}\|_{F}\sqrt{\sum_{j=1}^{m}\|\nabla\overline{\phi}_{j}^{U}\|_{L^{\infty}}^{2}}\leqslant 1.44\times 10^{-5}\times 0.025\times 1.33\leqslant 4.8\times 10^{-7}.$
@@ 4.5.2 Verification of Proposition 2
$M_{4}^{v}=(M_{5})^{2}-4M_{3}^{v}(\varepsilon^{v}+\varepsilon_{5})>0,$
where we recall
$M_{5}$ $:=$ $\frac{1}{4}-\eta_{1}-\overline{\lambda}-2K_{4}x^{U}_{1}-y^{U}_{1}-|\lambda^{\operatorname{con}}|-M_{2}^{v}.$
Step 1: Estimate of $\varepsilon_{5}$ and $|\lambda^{\operatorname{con}}|$.
Therefore, $\varepsilon_{5}$ and $\lambda^{\operatorname{con}}$ are bounded by
$|\lambda^{\operatorname{con}}|$ $\leqslant$ $\|\nabla U_{1}\|_{L^{2}}\|\overline{v}\|_{L^{\infty}}+\|\nabla U_{2}\|_{L^{\infty}}\leqslant 6.2\times 10^{-6}\times 2.52+4.8\times 10^{-7}\leqslant 1.62\times 10^{-5},$
$\varepsilon_{5}$ $\leqslant$ $(1.44\times 10^{-5}+1.3\times 10^{-7})\times 3.4+6.2\times 10^{-6}\times 2.52+4.8\times 10^{-7}\leqslant 6.6\times 10^{-5}.$
We have verified that the residuals satisfy
$\varepsilon^{v}+\varepsilon_{5}\leqslant 1.9\times 10^{-5}+6.6\times 10^{-5}\leqslant 8.5\times 10^{-5}.$
Step 2: Estimate of $M_{5}$.
Recall that $K_{4}\leqslant 0.091$ is introduced in Lemma 6. We can simplify $M_{5}$ as follows:
$M_{5}\geqslant\frac{1}{4}-0.199+0.113-2\times 0.091\times 6.2\times 10^{-6}-4.8\times 10^{-7}-1.62\times 10^{-5}-M_{2}^{v}\geqslant 0.1639-M_{2}^{v}.$
@@
Since $M_{1}^{v}$ is bounded by
$M_{1}^{v}=K^{v}_{5}+\frac{K_{6}^{v}}{(1-\varepsilon^{v}_{4})^{2}}\leqslant 0.47+\frac{0.46}{(1-5.3\times 10^{-5})^{2}}\leqslant 0.9301,$
we obtain the following upper bound for $M_{2}^{v}$:
$M_{2}^{v}\leqslant 0.005\times 0.9301+4.5\times 5.3\times 10^{-5}+0.004\leqslant 0.0089.$
$M_{5}$ thus has the following lower bound:
$M_{5}\geqslant 0.1639-0.0089=0.155.$
@@
$M_{0}$ $=$ $\|P_{\overline{v}}(\frac{1}{2}x\cdot\nabla\overline{v}-\Delta\overline{v}-\overline{U}\cdot\nabla\overline{v}+\overline{v}\cdot\nabla^{T}\overline{U})\|_{L^{2}}+\varepsilon_{5}$
$\leqslant$ $6.38+6.6\times 10^{-5}$
$\leqslant$ $6.4,$
which yields the following approximate values:
$M_{3}^{v}$ $\leqslant$ $6.4\times(1+1.02\times\sqrt{0.029})^{2}\leqslant 8.82,$
$M_{4}^{v}$ $\geqslant$ $0.155^{2}-4\times 8.82\times 8.5\times 10^{-5}\geqslant 0.021.$
@@
$x_{0}^{v}$ $=$ $\frac{2(\varepsilon^{v}+\varepsilon_{5})}{M_{5}+\sqrt{M_{4}^{v}}}\leqslant\frac{2\times 8.5\times 10^{-5}}{0.155+\sqrt{0.021}}\leqslant 5.67\times 10^{-4},$
$x_{1}^{v}$ $=$ $\sqrt{\frac{M_{5}}{1-2K_{4}\|\nabla U\|_{L^{2}}-M_{1}^{v}}}\frac{x_{0}^{v}}{2}\leqslant\sqrt{\frac{0.155}{1-2\times 0.091\times 6.2\times 10^{-6}-0.9301}}\times\frac{5.67\times 10^{-4}}{2}\leqslant 4.3\times 10^{-4}.$
@@
$|\lambda^{\operatorname{lin}}(v)|\leqslant M_{0}\|v\|_{L^{2}}\leqslant M_{0}(1+\sqrt{\sum_{j=1}^{m}\|G_{Q^{2}}^{v}\|\|\overline{\phi}_{j}^{v}\|_{L^{2}}^{2}})x_{0}^{v}\leqslant 6.4\times(1+0.18)\times 5.67\times 10^{-4}\leqslant 0.0044.$
@@
$|\lambda|\leqslant|\lambda^{\operatorname{lin}}(v)|+|\lambda^{\operatorname{con}}|\leqslant 0.0045\ll|\overline{\lambda}|\approx 0.113.$
This concludes the numerical verification and thus completes the proof of the proposition 2.